OpenStudy (anonymous):

tiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtiredtired

OpenStudy (jhannybean):

\[\large \sqrt[3]{x^2}*\sqrt[4]{x}= x^{2/3}*x^{1/4} = x^{(2/3)+(1/4)}\]

OpenStudy (jhannybean):

first simplify \[\large \frac{2}{3}+\frac{1}{4}\]

OpenStudy (jhannybean):

good job :) this is now the exponent of your base x.

OpenStudy (jhannybean):

 so your outer power is the power of your radical (or square root) and the inner power is the one that multiplies to your base "x"